\begin{thebibliography}{10}\itemsep=-1pt

\bibitem{bao2021beit}
Hangbo Bao, Li Dong, and Furu Wei.
\newblock Beit: Bert pre-training of image transformers.
\newblock {\em arXiv preprint arXiv:2106.08254}, 2021.

\bibitem{carion2020end}
Nicolas Carion, Francisco Massa, Gabriel Synnaeve, Nicolas Usunier, Alexander
  Kirillov, and Sergey Zagoruyko.
\newblock End-to-end object detection with transformers.
\newblock In {\em European conference on computer vision}, pages 213--229.
  Springer, 2020.

\bibitem{chen2022group}
Qiang Chen, Xiaokang Chen, Jian Wang, Haocheng Feng, Junyu Han, Errui Ding,
  Gang Zeng, and Jingdong Wang.
\newblock Group detr: Fast detr training with group-wise one-to-many
  assignment.
\newblock {\em arXiv preprint arXiv:2207.13085}, 2022.

\bibitem{chen2022context}
Xiaokang Chen, Mingyu Ding, Xiaodi Wang, Ying Xin, Shentong Mo, Yunhao Wang,
  Shumin Han, Ping Luo, Gang Zeng, and Jingdong Wang.
\newblock Context autoencoder for self-supervised representation learning.
\newblock {\em arXiv preprint arXiv:2202.03026}, 2022.

\bibitem{clark2020electra}
Kevin Clark, Minh-Thang Luong, Quoc~V Le, and Christopher~D Manning.
\newblock Electra: Pre-training text encoders as discriminators rather than
  generators.
\newblock {\em arXiv preprint arXiv:2003.10555}, 2020.

\bibitem{dai2021dynamic}
Xiyang Dai, Yinpeng Chen, Bin Xiao, Dongdong Chen, Mengchen Liu, Lu Yuan, and
  Lei Zhang.
\newblock Dynamic head: Unifying object detection heads with attentions.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 7373--7382, 2021.

\bibitem{deng2009imagenet}
Jia Deng, Wei Dong, Richard Socher, Li-Jia Li, Kai Li, and Li Fei-Fei.
\newblock Imagenet: A large-scale hierarchical image database.
\newblock In {\em 2009 IEEE conference on computer vision and pattern
  recognition}, pages 248--255. Ieee, 2009.

\bibitem{dosovitskiy2020image}
Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn,
  Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg
  Heigold, Sylvain Gelly, et~al.
\newblock An image is worth 16x16 words: Transformers for image recognition at
  scale.
\newblock {\em arXiv preprint arXiv:2010.11929}, 2020.

\bibitem{he2022masked}
Kaiming He, Xinlei Chen, Saining Xie, Yanghao Li, Piotr Doll{\'a}r, and Ross
  Girshick.
\newblock Masked autoencoders are scalable vision learners.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 16000--16009, 2022.

\bibitem{li2022dn}
Feng Li, Hao Zhang, Shilong Liu, Jian Guo, Lionel~M Ni, and Lei Zhang.
\newblock Dn-detr: Accelerate detr training by introducing query denoising.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 13619--13627, 2022.

\bibitem{li2022grounded}
Liunian~Harold Li, Pengchuan Zhang, Haotian Zhang, Jianwei Yang, Chunyuan Li,
  Yiwu Zhong, Lijuan Wang, Lu Yuan, Lei Zhang, Jenq-Neng Hwang, et~al.
\newblock Grounded language-image pre-training.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 10965--10975, 2022.

\bibitem{li2022exploring}
Yanghao Li, Hanzi Mao, Ross Girshick, and Kaiming He.
\newblock Exploring plain vision transformer backbones for object detection.
\newblock {\em arXiv preprint arXiv:2203.16527}, 2022.

\bibitem{lin2014microsoft}
Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva
  Ramanan, Piotr Doll{\'a}r, and C~Lawrence Zitnick.
\newblock Microsoft coco: Common objects in context.
\newblock In {\em European conference on computer vision}, pages 740--755.
  Springer, 2014.

\bibitem{liu2022dab}
Shilong Liu, Feng Li, Hao Zhang, Xiao Yang, Xianbiao Qi, Hang Su, Jun Zhu, and
  Lei Zhang.
\newblock Dab-detr: Dynamic anchor boxes are better queries for detr.
\newblock {\em arXiv preprint arXiv:2201.12329}, 2022.

\bibitem{liu2022swin}
Ze Liu, Han Hu, Yutong Lin, Zhuliang Yao, Zhenda Xie, Yixuan Wei, Jia Ning, Yue
  Cao, Zheng Zhang, Li Dong, et~al.
\newblock Swin transformer v2: Scaling up capacity and resolution.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 12009--12019, 2022.

\bibitem{liu2021swin}
Ze Liu, Yutong Lin, Yue Cao, Han Hu, Yixuan Wei, Zheng Zhang, Stephen Lin, and
  Baining Guo.
\newblock Swin transformer: Hierarchical vision transformer using shifted
  windows.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 10012--10022, 2021.

\bibitem{meng2021conditional}
Depu Meng, Xiaokang Chen, Zejia Fan, Gang Zeng, Houqiang Li, Yuhui Yuan, Lei
  Sun, and Jingdong Wang.
\newblock Conditional detr for fast training convergence.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 3651--3660, 2021.

\bibitem{peng2022beit}
Zhiliang Peng, Li Dong, Hangbo Bao, Qixiang Ye, and Furu Wei.
\newblock Beit v2: Masked image modeling with vector-quantized visual
  tokenizers.
\newblock {\em arXiv preprint arXiv:2208.06366}, 2022.

\bibitem{ridnik2021imagenet}
Tal Ridnik, Emanuel Ben-Baruch, Asaf Noy, and Lihi Zelnik-Manor.
\newblock Imagenet-21k pretraining for the masses.
\newblock {\em arXiv preprint arXiv:2104.10972}, 2021.

\bibitem{shao2019objects365}
Shuai Shao, Zeming Li, Tianyuan Zhang, Chao Peng, Gang Yu, Xiangyu Zhang, Jing
  Li, and Jian Sun.
\newblock Objects365: A large-scale, high-quality dataset for object detection.
\newblock In {\em Proceedings of the IEEE/CVF international conference on
  computer vision}, pages 8430--8439, 2019.

\bibitem{tarvainen2017mean}
Antti Tarvainen and Harri Valpola.
\newblock Mean teachers are better role models: Weight-averaged consistency
  targets improve semi-supervised deep learning results.
\newblock {\em Advances in neural information processing systems}, 30, 2017.

\bibitem{wang2022image}
Wenhui Wang, Hangbo Bao, Li Dong, Johan Bjorck, Zhiliang Peng, Qiang Liu, Kriti
  Aggarwal, Owais~Khan Mohammed, Saksham Singhal, Subhojit Som, et~al.
\newblock Image as a foreign language: Beit pretraining for all vision and
  vision-language tasks.
\newblock {\em arXiv preprint arXiv:2208.10442}, 2022.

\bibitem{wei2022contrastive}
Yixuan Wei, Han Hu, Zhenda Xie, Zheng Zhang, Yue Cao, Jianmin Bao, Dong Chen,
  and Baining Guo.
\newblock Contrastive learning rivals masked image modeling in fine-tuning via
  feature distillation.
\newblock {\em arXiv preprint arXiv:2205.14141}, 2022.

\bibitem{xie2022simmim}
Zhenda Xie, Zheng Zhang, Yue Cao, Yutong Lin, Jianmin Bao, Zhuliang Yao, Qi
  Dai, and Han Hu.
\newblock Simmim: A simple framework for masked image modeling.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 9653--9663, 2022.

\bibitem{xu2021end}
Mengde Xu, Zheng Zhang, Han Hu, Jianfeng Wang, Lijuan Wang, Fangyun Wei, Xiang
  Bai, and Zicheng Liu.
\newblock End-to-end semi-supervised object detection with soft teacher.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pages 3060--3069, 2021.

\bibitem{yang2022focal}
Jianwei Yang, Chunyuan Li, Xiyang Dai, and Jianfeng Gao.
\newblock Focal modulation networks, 2022.

\bibitem{zhai2022scaling}
Xiaohua Zhai, Alexander Kolesnikov, Neil Houlsby, and Lucas Beyer.
\newblock Scaling vision transformers.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and
  Pattern Recognition}, pages 12104--12113, 2022.

\bibitem{zhang2022dino}
Hao Zhang, Feng Li, Shilong Liu, Lei Zhang, Hang Su, Jun Zhu, Lionel~M Ni, and
  Heung-Yeung Shum.
\newblock Dino: Detr with improved denoising anchor boxes for end-to-end object
  detection.
\newblock {\em arXiv preprint arXiv:2203.03605}, 2022.

\bibitem{zhang2022glipv2}
Haotian Zhang, Pengchuan Zhang, Xiaowei Hu, Yen-Chun Chen, Liunian~Harold Li,
  Xiyang Dai, Lijuan Wang, Lu Yuan, Jenq-Neng Hwang, and Jianfeng Gao.
\newblock Glipv2: Unifying localization and vision-language understanding.
\newblock {\em arXiv preprint arXiv:2206.05836}, 2022.

\bibitem{zhu2020deformable}
Xizhou Zhu, Weijie Su, Lewei Lu, Bin Li, Xiaogang Wang, and Jifeng Dai.
\newblock Deformable detr: Deformable transformers for end-to-end object
  detection.
\newblock {\em arXiv preprint arXiv:2010.04159}, 2020.

\end{thebibliography}

\documentclass[10pt,twocolumn,letterpaper]{article}

\usepackage[pagenumbers]{cvpr}      

\usepackage{bbding}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{booktabs}

\usepackage[pagebackref,breaklinks,colorlinks]{hyperref}

\usepackage[capitalize]{cleveref}
\crefname{section}{Sec.}{Secs.}
\Crefname{section}{Section}{Sections}
\Crefname{table}{Table}{Tables}
\crefname{table}{Tab.}{Tabs.}

\def\cvprPaperID{*****} 
\def\confName{CVPR}
\def\confYear{2023}

\begin{document}


\title{Group DETR v2: Strong Object Detector with Encoder-Decoder Pretraining}
\author{
Qiang Chen$^{1*}$, Jian Wang$^{1}$\thanks{Equal contribution.}~~, Chuchu Han$^{1*}$, Shan Zhang$^2$, Zexian Li$^3$, Xiaokang Chen$^4$, Jiahui Chen$^3$ \\ Xiaodi Wang$^1$, Shuming Han$^1$, Gang Zhang$^1$, Haocheng Feng$^1$, Kun Yao$^1$, Junyu Han$^1$, Errui Ding$^1$\\ Jingdong Wang$^1$\thanks{Corresponding author.} \\
\\
$^1$Baidu VIS~~~~$^2$Australian National University~~~~$^3$Beihang University~~~~$^4$Peking University\\
}

\maketitle

\begin{abstract} 
We present a strong object detector with 
encoder-decoder pretraining and finetuning. 
Our method, called Group DETR v2, is built upon a vision transformer encoder ViT-Huge~\cite{dosovitskiy2020image}, 
a DETR variant DINO~\cite{zhang2022dino}, and an efficient DETR training method 
Group DETR~\cite{chen2022group}. 
The training process consists of self-supervised pretraining and finetuning a ViT-Huge encoder on ImageNet-1K, 
pretraining the detector on Object365, 
and finally finetuning it on COCO. 
Group DETR v2 achieves $\textbf{64.5}$ mAP on COCO test-dev, and
establishes a new SoTA on the COCO leaderboard\footnote{\url{https://paperswithcode.com/sota/object-detection-on-coco}}.
\end{abstract}

\section{Introduction} \label{sec:intro}

Recent studies show the effectiveness of transformer models at scale~\cite{dosovitskiy2020image,zhai2022scaling,liu2022swin}.
With encoder pretraining on large-scale data~\cite{deng2009imagenet,ridnik2021imagenet}, the models~\cite{bao2021beit,he2022masked,xie2022simmim,chen2022context,peng2022beit,wang2022image} are able to achieve superior results 
on various vision tasks, including object detection. 
With supervised encoder-decoder pretraining on a large-scale dataset, Object365~\cite{shao2019objects365}, DINO~\cite{zhang2022dino} achieves a state-of-the-art result on COCO~\cite{lin2014microsoft}.

Our method, {\em Group DETR v2}, is built upon ViT-Huge, DINO, and Group DETR. 
We adopt an encoder-decoder pretraining and finetuning pipeline: 
pretraining and then
finetuning a ViT-Huge encoder on ImageNet-1K~\cite{deng2009imagenet}, 
pretraining the detector, both the encoder and the decoder, on Object365, 
and finally finetuning it on COCO. 
Group DETR v2 achieves $\textbf{64.5}$ mAP on COCO test-dev~\cite{lin2014microsoft} (Table~\ref{tab:leaderboard} and Table~\ref{tab:coco_dev}), setting a new record for COCO object detection. 
We expect that the results will be further improved with more data and larger models. 

\section{Method} \label{sec:method}

\subsection{Architecture} \label{modeling}

\paragraph{Encoder.}
We adopt a ViT-Huge as the encoder.  We resort to the self-supervised pretrained model, Vit-Huge, e.g., with the MIM method CAE~\cite{chen2022context},
which shows superior performance on downstream tasks. 
We follow ViTDet~\cite{li2022exploring} to build multi-scale feature maps for multi-scale DETR.

\paragraph{Decoder.}
We adopt the transformer encoder-decoder framework as the decoder that shows promising detection results, including DETR~\cite{carion2020end}, Conditional DETR~\cite{meng2021conditional}, DAB-DETR~\cite{liu2022dab}, Deformable DETR~\cite{zhu2020deformable}, DN-DETR~\cite{li2022dn}, and DINO\cite{zhang2022dino}. Group DETR~\cite{chen2022group} provides further progress in improving the training convergence speed and the detection performance of various DETR variants. 
We build our detection decoder by combining DINO~\cite{zhang2022dino} 
into Group DETR~\cite{chen2022group}.

\begin{table}
  \centering
    \setlength{\tabcolsep}{5pt}
    \renewcommand{\arraystretch}{1.3}
    \footnotesize    
  \caption{\textbf{Our method, Group DETR v2, establishes
  a new SoTA on the COCO {\em test-dev} leaderboard.} }
  \begin{tabular}{c|cccccc}
    \toprule
    Method &  mAP & AP$50$ &  AP$75$ & AP$_s$ & AP$_m$ & AP$_l$ \\
    \midrule
    Group DETR v2  & $\mathbf{64.5}$ & $81.8$ & $71.1$ & $48.4$ & $67.2$ & $77.1$ \\
    \bottomrule
  \end{tabular}
  \label{tab:coco_dev}
\end{table}

\subsection{Implementation} \label{details}
The training process includes three stages: 
(i) pretrain and finetune the ViT-Huge encoder on ImageNet-1K~\cite{deng2009imagenet}, 
(ii) pretrain the whole detector (encoder and decoder) on Object365~\cite{shao2019objects365}, 
and (iii) finetune the detector on COCO~\cite{lin2014microsoft}. When pretraining the detector on Object365, 
we follow DINO~\cite{zhang2022dino} to only leave the first $5$k out of $80$k validation images as the validation set 
and add the other images to the training set. 
We also use other schemes when training the detector on Object365 and COCO, 
such as enlarging the image size to $1.5\times$ when finetuning and adopting test time augmentation. 
In addition, we apply the exponential moving average (EMA) technique~\cite{tarvainen2017mean}, use $100$ DN queries~\cite{zhang2022dino}, and adopt $11$ groups with Group DETR~\cite{chen2022group} during detector pretraining and finetuning. When finetuning the detector on COCO, we find that applying learning rate decay~\cite{clark2020electra,bao2021beit,he2022masked,chen2022context} for the components of the detector (encoder and decoder) gives a $\sim$$0.9$ mAP gain on COCO.

\section{Experiments}

\paragraph{Results on Object365 $\textbf{5}$k val.} 
We pretrain Group DETR v2 for $24$ epochs with $64$ A100 GPUs on Object365. 
On the first $5$k validation set, our Group DETR v2 achieves $\textbf{55.6}$ mAP. Table~\ref{tab:o365} gives detailed results.

\paragraph{Results on the COCO {\em test-dev}.} 
We finetune the detector
(pretrained on Object365) 
on the COCO training set
for $20$ epochs with $32$ A100 GPUs. 
During testing, we adopt test time augmentation 
with various scales and their flipped counterparts,
and perform fusion on the query features\footnote{According to our experiments, the fusion on the query features builds a robust feature across different scales and gives $\sim$$0.8$ mAP improvement.} and the final predictions~\cite{zhang2022dino}. 
Our Group DETR v2 achieves $\textbf{64.5}$ mAP on COCO {\em test-dev}. 
Table~\ref{tab:coco_dev} provides detailed results. 

\paragraph{Comparisons with state-of-the-art results on the COCO {\em test-dev} leaderboard.} 
We report the previous SoTA results on the COCO leaderboard. Table~\ref{tab:leaderboard} shows that only pretraining the ViT-Huge encoder on ImageNet-$1$K, Group DETR v2 outperforms other methods with larger models ({e.g.}, BEIT-3~\cite{wang2022image} and SwinV2-G~\cite{liu2022swin}) and more training data, and sets a new record on COCO {\em test-dev}. 
We expect that the results will be further improved with more data and larger models.

\begin{table}
  \centering
    \setlength{\tabcolsep}{5pt}
    \renewcommand{\arraystretch}{1.3}
    \footnotesize    
  \caption{\textbf{Results on Object365 $\textbf{5}$k val} with a single scale of $800\times 1333$.}
  \begin{tabular}{c|cccccc}
    \toprule
    Method &  mAP & AP$50$ &  AP$75$ & AP$_s$ & AP$_m$ & AP$_l$ \\
    \midrule
    Group DETR v2  & $\mathbf{55.6}$ & $68.8$ & $60.9$ & $36.6$ & $57.5$ & $71.3$ \\
    \bottomrule
  \end{tabular}
  \label{tab:o365}
\end{table}

{\small
\bibliographystyle{ieee_fullname}
\bibliography{egbib}
}

\end{document}